# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@ Abstract
Photonics represents one of the most promising approaches to large-scale quantum computation with millions of qubits and billions of gates, owing to the potential for room-temperature operation, high clock speeds, miniaturization of photonic circuits, and repeatable fabrication processes in commercial photonic foundries. We present an end-to-end architecture for fault-tolerant continuous-variable (CV) quantum computation using only passive on-chip components that can produce photonic qubits above the fault tolerance threshold with probabilities above 90%, and encodes logical qubits using physical qubits sampled from a distribution around the fault tolerance threshold. By requiring only low photon number resolution, the architecture enables the use of high-bandwidth photodetectors in CV quantum computing. Simulations of our qubit generation and logical encoding processes show a Gaussian cluster squeezing threshold of 12 dB to 13 dB. Additionally, we present a novel magic state generation protocol which requires only 13 dB of cluster squeezing to produce magic states with an order of magnitude higher probability than existing approaches, opening up the path to universal fault-tolerant quantum computation at less than 13 dB of cluster squeezing [^1] Approved for Public Release: Distribution Unlimited.
I Introduction
@@
Quantum computation promises the ability to exponentially speed up certain problems intractable to classical computers including simulations of large quantum systems, an observation first formalized by Richard Feynman Feynman (1982). Noisy intermediate scale quantum computers Preskill (2018) with $\approx 100$ physical qubits and $\approx 100$ to $1000$ quantum gates have been demonstrated on a variety of physical platforms Debnath et al. (2016); Kandala et al. (2017); Henriet et al. (2020); O’Malley et al. (2016), but it is widely expected that the computers must scale to millions of physical qubits and billions of operations to be useful for commercially valuable quantum simulations Kim et al. (2022); Gidney and Ekerå (2021); Litinski (2023). Matter based approaches face hurdles to get to such a computational volume, owing to the hardware requirements of low temperatures, slow gate rates, or lossy conversions between matter and photons Meesala et al. (2024); Kandala et al. (2017); O’Malley et al. (2016); Henriet et al. (2020); Saffman (2016). Photonics provides a promising path towards large-scale quantum computing with scalable generation and manipulation of qubits at room temperature, high-speed measurements, miniaturization using photonic integrated circuits, and repeatable low-loss mass manufacturing at commercial foundries.
@@
Photonic quantum information can be encoded into either discrete quantum degrees of freedom of individual single photons, known as the discrete variable (DV) approach Kok et al. (2007); Bartolucci et al. (2021), or in continuous-valued orthogonal quadratures of the electric field, known as the continuous variable (CV) approach Lloyd and Braunstein (1999); Fabre and Treps (2020). CV quantum photonics brings an advantage over DV approaches, namely the deterministic generation of entanglement over large photonic quantum states Weedbrook et al. (2012); Menicucci (2014). Millions of entangled optical modes have been experimentally generated in CV systems Asavanant et al. (2019). This makes CV photonics ideal for measurement-based QC (MBQC) implementations Menicucci et al. (2006), in which quantum computation is performed via single-qubit measurements on a large entangled quantum state. In photonics, single modes or collections of modes in the entangled state act as the analogues of qubits in matter-based approaches.
@@
The key concept to enable quantum computation with large-scale entangled states, or in general with a large number of qubits, is that of fault tolerance. Fault-tolerant quantum computing (FTQC) is an approach to QC in which a computational (logical) qubit is encoded in many physical qubits on which quantum error correction (QEC) protocols are employed to correct for errors such as decoherence and photon loss Gottesman et al. (2001); Cai et al. (2021); Xu et al. (2023); If the probability of physical errors is below a certain value determined by the QEC protocol, called the fault tolerance threshold, the probability of logical error can be exponentially suppressed. There exist no direct fault-tolerant protocols for quadrature encoded continuous-variable quantum computing (CVQC) Ohliger (2010). An alternative approach is to discretise the CV space into qubits Ohliger (2010); Menicucci (2014).
@@
The most promising qubits for CV FTQC are the Gottesman-Kitaev-Preskill (GKP) qubit Gottesman et al. (2001) and the magic states Bravyi and Kitaev (2005). Native logical gates are implemented with Gaussian operations on GKP qubits, while universal quantum computation is possible with the ability to create magic states. In fact, using only CV Gaussian resources combined with GKP states, one has all the necessary ingredients to achieve universal CVQC Baragiola et al. (2019). Therefore, much effort in CVQC has gone into devising scalable methods of producing high quality GKP qubits (that is, with high effective squeezing; see Sec. II), but producing these states with high success rates in photonics has remained a challenge (see Sec. V). After obtaining high quality GKP and magic qubits, CV architectures achieve fault tolerance by concatenating a second qubit-level error correcting code with the GKP states as the physical qubits. To implement these codes with optical fields, photonic switches were proposed either to increase the low success probabilities of GKP state generation Bourassa et al. (2021); Tzitrin et al. (2021) or for quantum error correction Larsen et al. (2021) (see Sec. VIII). Great progress has been made in lowering the noise and losses in photonic switches, but the loss levels are two orders-of-magnitude higher than passive photonics Team (2024), making removal of switches from the architecture a desirable goal.
@@
With this context, we present four contributions in synthesizing an end-to-end CV FTQC architecture. First, we provide a protocol that produces fault-tolerant CV photonic qubits –GKP qubits– at state-of-the-art probabilities of more than $90\%$. Notably, the architecture makes use of all physical qubits, both above and below the FTQC threshold, to construct logical qubits. Second, we provide methods to generate GKP qubits on an integrated switchless photonic chip, opening up the possibility of low-loss manufacturability of the photonic chips with passive components. Our methods for state generation require only low photon number resolution (PNR) for which either superconducting microstrip single photon detectors (SMSPDs) Kong et al. (2024) or room temperature photon number resolving detectors Nehra et al. (2020) would suffice. Third, we present protocols for magic state generation with state-of-the-art success probability, an order of magnitude better than existing methods, for universal FTQC with feasible resource requirements. Fourth, we present the results of our quantum error correction simulations including GKP resource generation costs. We arrive at the gaussian squeezing threshold, the quantum noise metric in CVQC Menicucci (2014); Schnabel (2017), of $12$ dB to $13$ dB for fault tolerance, and distillable magic state success probability of $4.8\%$ also at $13$ dB. To our knowledge, the presented simulations are the first end-to-end CVQC simulations incorporating a realistic process producing GKP states with a range of effective squeezing values for fault-tolerant operation.
@@
II CVQC Toolkit and Non-Gaussian States
@@
In photonic CV MBQC, the set of entangling operations, non-destructive gates, and destructive measurements come from the CV photonics toolkit Pfister (2019), the most prominent components of which are the squeezing operations, linear optics, quadrature displacements, homodyne measurements, and photon number resolving measurements. In the following, we define the key terms, including entangled states, measurement operations, and non-Gaussian modes, which form the basic building blocks of our architecture.
@@
The entangled state often used in photonic CV MBQC is the cluster state, canonically defined as Gaussian-squeezed modes entangled using controlled-phase gates ($C_{Z}$ielsen (2006). While CVQC was initially investigated with Gaussian cluster states requiring inline squeezing due to the use of controlled-phase gates Raussendorf and Briegel (2001); Flammia et al. (2009), an alternative formalism exists where inline squeezing is replaced by offline squeezing and passive interferometers Menicucci (2011); Walshe et al. (2021), resulting in a macronode cluster state, where collections of physical optical modes can be treated as a single mode on an underlying lattice Alexander and Menicucci (2016). An example of macronode cluster is the multimode state called a dual rail quantum wire, generated using the two-mode-squeezing operation implemented in a non-linear crystal followed by passive beamsplitters Pysher et al. (2009).
@@
The name squeezing stems from the shape of the Wigner quasiprobability distribution of the conjugate photonic quadratures (denoted $p$ and $q$), in which the distribution for one quadrature has a narrower width compared to a coherent state, while the other quadrature has a broader distribution. The term also refers to the quantum noise on a measurement along one or a combination of the quadrature axes in phase space. In CV MBQC, because all gates are carried out via measurement operations, the noise on these gates is parametrized by the squeezing observed in each measurement. Thus, physical gate noise is directly tied to the squeezing level associated with each mode being measured. Ideal Gaussian squeezed states (infinitely squeezed) are denoted by |$\ket{m}_{p}$, where $m$ is the eigenvalue of the quadrature $p$. For the results presented in this paper, all simulations are run with finitely squeezed states (i.e., resulting from applying the finite squeezing (single-mode or two-mode) operator to vacuum). FT thresholds in terms of squeezing are important for CV FTQC as single and two-mode squeezing forms the fundamental resource for CVQC Gu et al. (2009); Menicucci et al. (2007).
@@
The operations used to project a quantum state onto quadratures eigenstates are called homodyne measurements Lvovsky and Raymer (2009): a homodyne measurement with outcome $m$ on quadrature $p$ projects the measured mode onto the state |$\ket{m}_{p}$. An important property of homodyne measurements chosen with appropriate bases is that they can be used to either delete a node from a cluster state or teleport the quantum information from one node to another Gu et al. (2009), allowing one to shape cluster states during quantum computation.
@@
While deterministic and scalable Gaussian entanglement generation gives us a substrate quantum state to use in CVQC, and homodyne measurements can be used to shape the cluster state and implement quantum gates, universal CVQC requires additional resources to enable quantum computations that are not efficiently classically simulable Bartlett et al. (2002). One prominent operation to enable universal CVQC is photon subtraction Ourjoumtsev et al. (2006), which can be used to implement gates consisting of arbitrary powers of the quadrature operators on squeezed states. Experimentally, in quantum optics, the most common method for performing photon subtraction involves using a highly reflective beamsplitter, with a photon detector—capable or not of resolving the number of detected photons— placed in the transmission arm Ourjoumtsev et al. (2006); Melalkia et al. (2022). Throughout this article, whenever the photon subtraction process is mentioned, we refer to this experimental setup. One of the key applications of these higher-order gates is the production of non-Gaussian states in an entangled cluster, which along with homodyne detections and displacements, give us the ability to apply quantum error correction and operate the universal quantum computer above the FT threshold Menicucci (2014).
@@
The fundamental non-Gaussian resource states in our architecture are the squeezed cat states |±$\ket{\pm\alpha,r}$, which are superpositions of displaced squeezed states, i.e., |±$\ket{\pm\alpha,r}\sim(D(\alpha)+D(-\alpha))\mathcal{S}(r)\ket{0}$, where $D$ is the displacement operator in the phase space, $D(\alpha)=\exp{(\alpha a^{\dagger}-\alpha^{\star}a)}$, $\mathcal{S}$ is the squeezing operator, $\mathcal{S}(r)=\exp(\frac{1}{2}(r^{*}a^{2}-ra^{\dagger^{2}}))$, and $a^{\dagger}$ and $a$ are the creation and annihilation operators, respectively. Real $\alpha$ results in $q$ quadrature displacements ($D_{q}$), while imaginary values result in $p$ displacements ($D_{p}$). Cat states are parametrized by squeezing ($r$) and displacement ($\alpha$), which we take to be real-valued in the following results. The Wigner function for a squeezed cat state is shown in Fig. 1. While FTQC codes exist with cat states as the logical qubits, these require the experimentally-challenging parity check operator Schlegel et al. (2022), limiting the role of cat states in photonic FTQC to that of resource states which we consume to produce qubits.
@@
GKP states, and more generally grid states, were first proposed in [Ref. Gottesman et al. (2001)], with a considerable amount of effort spent studying their use in fault-tolerant architectures in recent years Tzitrin et al. (2020); Mensen et al. (2021); Grimsmo and Puri (2021); a typical Wigner function is shown in 1. The sensor state, denoted |$\ket{S}$, is a symmetric grid state defined as the $+1$ eigenstate of the displacements $D_{q}(\sqrt{2\pi})$ and $D_{p}(\sqrt{2\pi})$, while a GKP qubit is the $+1$ eigenstate of the displacements $D_{q}(2\sqrt{\pi})$ and $D_{p}(\sqrt{\pi})$, where we have set $[a,a^{\dagger}]=1$  Gottesman et al. (2001); Grimsmo and Puri (2021) (see Appendix Sec. XI.1 for more details). Interference between sensor states using passive linear optics can be used to produce entangled GKP Bell pairs  Walshe et al. (2020).
@@
Ideal GKP states have wavefunctions that are a series of delta functions in the $q$-basis (or $p$-basis) of the continuous-variable $(q,p)$ phase space of a bosonic mode, and are therefore unphysical (see Appendix XI.1 for ideal GKP wavefunctions). Physically-relevant approximations to the ideal GKP states can be obtained by applying the damping operator to the ideal state Matsuura et al. (2020). The width of the peaks provides an effective squeezing parameter, which describes the closeness of a given GKP state to the ideal state. The probability of physical gate errors in CVQC depends on the effective squeezing of the underlying GKP states, which denotes the noise level of measurements associated with the qubit compared to the standard noise of a coherent state Weigand and Terhal (2018); minimizing this noise corresponds to reduced physical error rates. For any given density matrix $\rho$ and the desired displacement operator with displacement $u$, we estimate the effective squeezing of the output state as $\Delta_{D}=\frac{1}{u}\sqrt{\ln{|TrD(u)\rho|}}$ where $Tr$ is the trace operation Weigand and Terhal (2018). The quality of a realistic GKP state for FTQC is quantified most readily by the effective squeezing parameter since the probability of error in the logical X or Z operations is bounded by $P<\frac{2\Delta}{\pi}\exp{-\pi/(4\Delta^{2})}$, where $\Delta$ is the effective squeezing in the respective quadrature Weigand and Terhal (2018). Note that the GKP state can have different effective squeezing in the two quadratures owing to different parameters at play in the generation protocols. A final important non-Gaussian state in our architecture is the magic state. These are essential for universal quantum computation by providing an implementation of a complete gate set. In CVQC, magic states are also grid states; their Wigner functions contain periodic structure (see Fig. 1; also Appendix Sec. XI.1), and they can be derived from CV GKP states Baragiola et al. (2019). An important metric is the distillation threshold of magic states, which quantifies the fidelity above which many copies of noisy magic states can be used to distill a high fidelity magic state useful for applying fault-tolerant non-Clifford gates, e.g., the $T$ gate Bravyi and Kitaev (2005).
@@
III Architecture and Design Overview
Our photonic QC architecture contains three key components for physical qubit generation (see Fig 2): cat state generation via the Photon-counting-Assisted Node-Teleportation Method (PhANTM) Eaton et al. (2022a) (see also section IV), an adaptive breeding protocol (detailed in section V), and a magic state generation protocol (described in section VI). The logical layer consists of two major components: logical qubit generation, which consists of distributing physical qubits as the nodes within the Raussendorf-Harrington-Goyal (RHG) lattice Raussendorf et al. (2006); and the quantum error correction (QEC) encoding/decoding system, which is also responsible for running logical instructions. All of the physical qubit generation and RHG lattice generation components are possible on a set of integrated photonic chips without active photonic switches or variable beamsplitters on the quantum modes to minimize loss and noise. While fast electro-optic modulation is needed on local oscillators during homodyne measurements, the local oscillator power can be corrected to account for any losses.

### v3

@@ Abstract
Photonics presents one of the most promising approaches to large-scale quantum computation with millions of qubits and billions of gates, owing to the potential for room-temperature operation, high clock speeds, miniaturization of photonic circuits, and repeatable fabrication processes in commercial photonic foundries. We present an end-to-end architecture for fault-tolerant continuous-variable (CV) quantum computation using only passive on-chip components that can produce photonic qubits above the fault tolerance threshold with probabilities above 90%, and encodes logical qubits using physical qubits sampled from a distribution around the fault tolerance threshold. By requiring only low photon number resolution, the architecture enables the use of high-bandwidth photodetectors in CV quantum computing. Simulations of our qubit generation and logical encoding processes show a Gaussian cluster squeezing threshold of 12 dB to 13 dB. Additionally, we present a novel magic state generation protocol which requires only 13 dB of cluster squeezing to produce magic states with an order of magnitude higher probability than existing approaches, opening up the path to universal fault-tolerant quantum computation at less than 13 dB of cluster squeezing.
[…]
1 Introduction
@@
Quantum computation promises the ability to exponentially speed up certain problems intractable to classical computers including simulations of large quantum systems, an observation first formalized by Richard Feynman [1]. Noisy intermediate scale quantum computers [2] with $\approx 100$ physical qubits and $\approx 100$ to $1000$ quantum gates have been demonstrated on a variety of physical platforms [3, 4, 5, 6], but it is widely expected that the computers must scale to millions of physical qubits and billions of operations to be useful for commercially valuable quantum simulations [7, 8, 9]. Matter based approaches face hurdles to get to such a computational volume, owing to the hardware requirements of low temperatures, slow gate rates, or lossy conversions between matter and photons [10, 4, 6, 5, 11]. Photonics provides a promising path towards large-scale quantum computing with scalable generation and manipulation of qubits at room temperature, high-speed measurements, miniaturization using photonic integrated circuits, and repeatable low-loss mass manufacturing at commercial foundries.
@@
Photonic quantum information can be encoded into either discrete quantum degrees of freedom of individual single photons, known as the discrete variable (DV) approach [12, 13], or in continuous-valued orthogonal quadratures of the electric field, known as the continuous variable (CV) approach [14, 15]. CV quantum photonics brings an advantage over DV approaches, namely the deterministic generation of entanglement over large photonic quantum states [16, 17]. Millions of entangled optical modes have been experimentally generated in CV systems [18]. This makes CV photonics ideal for measurement-based QC (MBQC) implementations [19], in which quantum computation is performed via single-qubit measurements on a large entangled quantum state. In photonics, single modes or collections of modes in the entangled state act as the analogues of qubits in matter-based approaches.
@@
The key concept to enable quantum computation with large-scale entangled states, or in general with a large number of qubits, is that of fault tolerance. Fault-tolerant quantum computing (FTQC) is an approach to QC in which a computational (logical) qubit is encoded in many physical qubits on which quantum error correction (QEC) protocols are employed to correct for errors such as decoherence and photon loss [20, 21, 22]; If the probability of physical errors is below a certain value determined by the QEC protocol, called the fault tolerance threshold, the probability of logical error can be exponentially suppressed. There exist no direct fault-tolerant protocols for quadrature encoded continuous-variable quantum computing (CVQC) [23]. An alternative approach is to discretise the CV space into qubits [23, 17].
@@
The most promising qubits for CV FTQC are the Gottesman-Kitaev-Preskill (GKP) qubit [20] and the magic states [24]. Native logical gates are implemented with Gaussian operations on GKP qubits, while universal quantum computation is possible with the ability to create magic states. In fact, using only CV Gaussian resources combined with GKP states, one has all the necessary ingredients to achieve universal CVQC [25]. Therefore, much effort in CVQC has gone into devising scalable methods of producing high quality GKP qubits (that is, with high effective squeezing; see Sec. 2), but producing these states with high success rates in photonics has remained a challenge (see Sec. 5). After obtaining high quality GKP and magic qubits, CV architectures achieve fault tolerance by concatenating a second qubit-level error correcting code with the GKP states as the physical qubits. To implement these codes with optical fields, photonic switches were proposed either to increase the low success probabilities of GKP state generation [26, 27] or for quantum error correction [28] (see Sec. 8). Great progress has been made in lowering the noise and losses in photonic switches, but the loss levels are two orders-of-magnitude higher than passive photonics [29], making removal of switches from the architecture a desirable goal.
@@
With this context, we present four contributions in synthesizing an end-to-end CV FTQC architecture. First, we provide a protocol that produces fault-tolerant CV photonic qubits –GKP qubits– at state-of-the-art probabilities of more than $90\%$. Notably, the architecture makes use of all physical qubits, both above and below the FTQC threshold, to construct logical qubits. Second, we provide methods to generate GKP qubits on an integrated switchless photonic chip, opening up the possibility of low-loss manufacturability of the photonic chips with passive components. Our methods for state generation require only low photon number resolution (PNR) for which either superconducting microstrip single photon detectors (SMSPDs) [30] or room temperature photon number resolving detectors [31] would suffice. Third, we present protocols for magic state generation with state-of-the-art success probability, an order of magnitude better than existing methods, for universal FTQC with feasible resource requirements. Fourth, we present the results of our quantum error correction simulations including GKP resource generation costs. We arrive at the gaussian squeezing threshold, the quantum noise metric in CVQC [17, 32], of $12$ dB to $13$ dB for fault tolerance, and distillable magic state success probability of $4.8\%$ also at $13$ dB. To our knowledge, the presented simulations are the first end-to-end CVQC simulations incorporating a realistic process producing GKP states with a range of effective squeezing values for fault-tolerant operation.
@@
2 CVQC Toolkit and Non-Gaussian States
@@
In photonic CV MBQC, the set of entangling operations, non-destructive gates, and destructive measurements come from the CV photonics toolkit [33], the most prominent components of which are the squeezing operations, linear optics, quadrature displacements, homodyne measurements, and photon number resolving measurements. In the following, we define the key terms, including entangled states, measurement operations, and non-Gaussian modes, which form the basic building blocks of our architecture.
@@
The entangled state often used in photonic CV MBQC is the cluster state, canonically defined as Gaussian-squeezed modes entangled using controlled-phase gates ($C_{Z}$[34]. While CVQC was initially investigated with Gaussian cluster states requiring inline squeezing due to the use of controlled-phase gates [35, 36, 19], an alternative formalism exists where inline squeezing is replaced by offline squeezing and passive interferometers [37, 38, 39], resulting in a macronode cluster state, where collections of physical optical modes can be treated as a single mode on an underlying lattice [40]. An example of macronode cluster is the multimode state called a dual rail quantum wire, generated using the two-mode-squeezing operation implemented in a non-linear crystal followed by passive beamsplitters [41].
@@
The name squeezing stems from the shape of the Wigner quasiprobability distribution of the conjugate photonic quadratures (denoted $p$ and $q$), in which the distribution for one quadrature has a narrower width compared to a coherent state, while the other quadrature has a broader distribution. The term also refers to the quantum noise on a measurement along one or a combination of the quadrature axes in phase space. In CV MBQC, because all gates are carried out via measurement operations, the noise on these gates is parametrized by the squeezing observed in each measurement. Thus, physical gate noise is directly tied to the squeezing level associated with each mode being measured. Ideal Gaussian squeezed states (infinitely squeezed) are denoted by |$\ket{m}_{p}$, where $m$ is the eigenvalue of the quadrature $p$. For the results presented in this paper, all simulations are run with finitely squeezed states (i.e., resulting from applying the finite squeezing (single-mode or two-mode) operator to vacuum). FT thresholds in terms of squeezing are important for CV FTQC as single and two-mode squeezing forms the fundamental resource for CVQC [42, 43].
@@
The operations used to project a quantum state onto quadratures eigenstates are called homodyne measurements [44]: a homodyne measurement with outcome $m$ on quadrature $p$ projects the measured mode onto the state |$\ket{m}_{p}$. An important property of homodyne measurements chosen with appropriate bases is that they can be used to either delete a node from a cluster state or teleport the quantum information from one node to another [42], allowing one to shape cluster states during quantum computation.
@@
While deterministic and scalable Gaussian entanglement generation gives us a substrate quantum state to use in CVQC, and homodyne measurements can be used to shape the cluster state and implement quantum gates, universal CVQC requires additional resources to enable quantum computations that are not efficiently classically simulable [45]. One prominent operation to enable universal CVQC is photon subtraction [46], which can be used to implement gates consisting of arbitrary powers of the quadrature operators on squeezed states. Experimentally, in quantum optics, the most common method for performing photon subtraction involves using a highly reflective beamsplitter, with a photon detector—capable or not of resolving the number of detected photons— placed in the transmission arm [46, 47]. Throughout this article, whenever the photon subtraction process is mentioned, we refer to this experimental setup. One of the key applications of these higher-order gates is the production of non-Gaussian states in an entangled cluster, which along with homodyne detections and displacements, give us the ability to apply quantum error correction and operate the universal quantum computer above the FT threshold [17].
@@
The fundamental non-Gaussian resource states in our architecture are the squeezed cat states |±$\ket{\pm\alpha,r}$, which are superpositions of displaced squeezed states, i.e., |±$\ket{\pm\alpha,r}\sim(D(\alpha)+D(-\alpha))\mathcal{S}(r)\ket{0}$, where $D$ is the displacement operator in the phase space, $D(\alpha)=\exp{(\alpha a^{\dagger}-\alpha^{\star}a)}$, $\mathcal{S}$ is the squeezing operator, $\mathcal{S}(r)=\exp(\frac{1}{2}(r^{*}a^{2}-ra^{\dagger^{2}}))$, and $a^{\dagger}$ and $a$ are the creation and annihilation operators, respectively. Real $\alpha$ results in $q$ quadrature displacements ($D_{q}$), while imaginary values result in $p$ displacements ($D_{p}$). Cat states are parametrized by squeezing ($r$) and displacement ($\alpha$), which we take to be real-valued in the following results. The Wigner function for a squeezed cat state is shown in Fig. 1. While FTQC codes exist with cat states as the logical qubits, these require the experimentally-challenging parity check operator [48], limiting the role of cat states in photonic FTQC to that of resource states which we consume to produce qubits.
@@
GKP states, and more generally grid states, were first proposed in [Ref. [20]], with a considerable amount of effort spent studying their use in fault-tolerant architectures in recent years [49, 50, 51]; a typical Wigner function is shown in 1. The sensor state, denoted |$\ket{S}$, is a symmetric grid state defined as the $+1$ eigenstate of the displacements $D_{q}(\sqrt{2\pi})$ and $D_{p}(\sqrt{2\pi})$, while a GKP qubit is the $+1$ eigenstate of the displacements $D_{q}(2\sqrt{\pi})$ and $D_{p}(\sqrt{\pi})$, where we have set $[a,a^{\dagger}]=1$  [20, 51] (see Appendix Sec. 11.1 for more details). Interference between sensor states using passive linear optics can be used to produce entangled GKP Bell pairs  [52].
@@
Ideal GKP states have wavefunctions that are a series of delta functions in the $q$-basis (or $p$-basis) of the continuous-variable $(q,p)$ phase space of a bosonic mode, and are therefore unphysical (see Appendix 11.1 for ideal GKP wavefunctions). Physically-relevant approximations to the ideal GKP states can be obtained by applying the damping operator to the ideal state [17, 53]. The width of the peaks provides an effective squeezing parameter, which describes the closeness of a given GKP state to the ideal state. The probability of physical gate errors in CVQC depends on the effective squeezing of the underlying GKP states, which denotes the noise level of measurements associated with the qubit compared to the standard noise of a coherent state [54]; minimizing this noise corresponds to reduced physical error rates. For any given density matrix $\rho$ and the desired displacement operator with displacement $u$, we estimate the effective squeezing of the output state as $\Delta_{D}=\frac{1}{u}\sqrt{\ln{|TrD(u)\rho|}}$ where $Tr$ is the trace operation [54]. The quality of a realistic GKP state for FTQC is quantified most readily by the effective squeezing parameter since the probability of error in the logical X or Z operations is bounded by $P<\frac{2\Delta}{\pi}\exp{-\pi/(4\Delta^{2})}$, where $\Delta$ is the effective squeezing in the respective quadrature [54]. Note that the GKP state can have different effective squeezing in the two quadratures owing to different parameters at play in the generation protocols. A final important non-Gaussian state in our architecture is the magic state. These are essential for universal quantum computation by providing an implementation of a complete gate set. In CVQC, magic states are also grid states; their Wigner functions contain periodic structure (see Fig. 1; also Appendix Sec. 11.1), and they can be derived from CV GKP states [25]. An important metric is the distillation threshold of magic states, which quantifies the fidelity above which many copies of noisy magic states can be used to distill a high fidelity magic state useful for applying fault-tolerant non-Clifford gates, e.g., the $T$ gate [24].
@@
3 Architecture and Design Overview
Our photonic QC architecture contains three key components for physical qubit generation (see Fig 2): cat state generation via the Photon-counting-Assisted Node-Teleportation Method (PhANTM) [55] (see also section 4), an adaptive breeding protocol (detailed in section 5), and a magic state generation protocol (described in section 6). The logical layer consists of two major components: logical qubit generation, which consists of distributing physical qubits as the nodes within the Raussendorf-Harrington-Goyal (RHG) lattice [56]; and the quantum error correction (QEC) encoding/decoding system, which is also responsible for running logical instructions. All of the physical qubit generation and RHG lattice generation components are possible on a set of integrated photonic chips without active photonic switches or variable beamsplitters on the quantum modes to minimize loss and noise. While fast electro-optic modulation is needed on local oscillators during homodyne measurements, the local oscillator power can be corrected to account for any losses.